OpenStudy (lilai3):

PLEASE HELP!!! IT WILL BE GREATLY APPRECIATED, AND I WILL FAN/REWARD MEDAL!!!! (; Raquel is mixing lemon-lime soda and a fruit juice blend that is 45% juice. If she uses 3 quarts of soda, how many quarts of fruit juice must be added to produce punch that is 30% juice? Thank you.

OpenStudy (azureilai):

Use proportions (3/x)=(70/100) find x and then multiply that answer by (3/10)

OpenStudy (lilai3):

 please use this chart to fill it in. the *** is the column that i'm not sure what category to fit that in

OpenStudy (azureilai):

I don't know what to put it either... I is that all the instructions for the answer?

OpenStudy (lilai3):

uhm, no, but my teacher told me to do it on a graph, so i made it.

OpenStudy (azureilai):

ah, I'm sorry, I really don't know then. Because this isn't really a graphable answer, there are no independent or dependent variables, it is just plain calculation. You might want to check with your classmates or teacher again. Sorry for not being able to help

OpenStudy (lilai3):

it's alright. @silversurfer ? anything?

OpenStudy (anonymous):

nope couldn't solve

OpenStudy (lilai3):

@Directrix ?

Directrix (directrix):

I don't know about this chart. That is, the purpose of the chart.

OpenStudy (lilai3):

well, my teacher said that the chart must be filled out, but i can't see anyway you could make a chart out of that...can you?

Directrix (directrix):

Maybe, we can work the problem and find the answer. After that, the chart may make sense.

OpenStudy (lilai3):

i'm sorry i don't get the answer

Directrix (directrix):

Let x = quarts of juice to be added .45x is the pure juice in the blend Total liquid in blend = (3 + x) quarts Pure Juice in blend: .3(3 + x) .3(3 + x) = .45x x = 6 so you are adding 6 quarts of pure juice to get a blend that is 30% pure juice. @lilai3 I don't know about the chart. If it had labels for the columns and rows, that would help.

Directrix (directrix):

@lilai3 I have an idea about this chart. Just a few moments more.

OpenStudy (lilai3):

okay thank you!

Directrix (directrix):

Here's the Chart

OpenStudy (lilai3):